Fermi National Laboratory


Accelerator Update - May 23-May 28

Friday May 23
The day shift began with the Tevatron (TeV) in shot setup, and with MiniBooNe taking beam.

At 12:27 PM, Operations established store 2600 with an initial luminosity of 20.4E30. The Antiproton Source (Pbar) resumed stacking.

A Booster power supply (B:V803) tripped off due to an overtemp indication at 9:42 PM. Operators reset the supply and stacking resumed at 10:07 PM.

At 11:59 PM, Operations terminated store 2600.

Saturday May 24
The midnight shift began with Pbar cooling the antiproton beam, and with MiniBooNe taking beam.

At 3 AM, a Linac RF station (KRF3) was tripping off regularly. Operators investigated and found that another station (LRF5) was causing the problems. Operators turned up the filament power, which wasn't good for this aging power amplifier tube, but there was little choice.

A Pbar expert discovered a quadrupole shunt SCR off in the AP3 line at 3:06 AM. The expert made system adjustments to compensate for the bad shunt. A technician will come in to repair this device later in the morning.

Operations rested LRF5 after loading the final TeV protons at 4:17 AM. This turns off the MiniBooNe beam.

At 4:47 AM, Operations established store 2604 with an initial luminosity of 34.41E30.

Operations turned LRF5 back on at 5:31 AM. Pbar resumed stacking and MiniBooNe resumed taking beam.

Operations stopped the Pbar stack, at 8:31 AM, so that techs could access the AP3 transfer line and repair the quadrupole shunt. Stacking resumed at 12:43 PM.

At 7:40 PM, Recycler experts took shots of protons from the Main Injector (MI) for studies.

A Booster RF station (BRF11) tripped off and wouldn't reset. Operators bypassed the station at 9:40 PM.

Operations terminated store 2604 at 10 PM and put the TeV back into shot setup at 10:53 PM.

At 11:39 PM, operators rested LRF5. No beam to MiniBooNe.

Sunday May 25
The midnight shift began with the TeV in shot setup, and with LRF5 allowed to rest after the protons were loaded.

At 12:12 AM, Operations established store 2606 with an initial luminosity of 31.43E30.

Operators woke up LRF5 at 12:52 AM. Pbar resumed stacking and MiniBooNe resumed taking beam.

At 1:22 AM, LRF3 tripped off and reset. Then at 1:30 AM, LRF2 tripped off and reset.

Operations turned off the Booster beam at 9:43 AM so techs could repair BRF11. The techs had the station disconnected from the system by 10:14 AM so they could begin repairs. Pbar and MiniBooNe resumed taking beam.

At 3:13 PM, the RF techs connected BRF11 back into the system.

BRF11 tripped off at 3:25 PM and again at 3:38 PM; it reset okay both times.

At 4 PM, Operations terminated store 2604.

The Pbar Lithium Lens had a water pressure alarm at 4:30 PM. Experts advised operators to raise the pressure limit from 13 to 14 psi. This was done at 5:25 PM.

Operations put the TeV into shot setup at 4:42 PM.

At 6:15 PM, Operations rested LRF5.

The TeV suffered a sector E1 quench at 6:34 PM. Operators believe this was a tune problem. Pbar stacking resumed at 7 PM.

At 8:16 PM, a TeV expert arrived to investigate the TeV quench. They never found a good cause for the quench.

Operations put the TeV through at dry squeeze at 10 PM.

Monday May 26
The midnight shift began with Operators conducting TeV studies, with Pbar stacking, and with MiniBooNe taking beam.

At 4:53 AM, Linac began suffering RF problems. Operators turned up the filament power on LRF5 and this helped. At 6:35 AM, KRF5 tripped off and reset, and then, at 7:21 AM, KRF6 tripped off and reset.

Operations put the TeV into shot setup at 8 AM.

The TeV protons finished loading at 9:53 AM and LRF5 was put down for a nap.

At 10:28 AM, Operations established store 2614 with an initial luminosity of 30.01E30.

A MI RF station (MIRF8) tripped off at 10:55 AM. Over the next three and a half hours, the station tripped off and was reset four more times.

At 11:11 AM, a Pbar Debuncher device (D:PMAGV) tripped off due to no water flow. Operation contacted the Pbar experts, and one came in to restart the system at noon. Stacking resumed at 1:22 PM, but there was still a temperature problem and possibly a small leak in the LCW system. The stacking rate was very poor. Near the end of the day shift Pbar experts reported that there appeared to be a short in the Lithium Lens transformer.

The MI power supply loop, located at MI-30, tripped off at 3:07 PM due to overcurrent. It was still tripping off at 3:24 PM, so an expert was called in to investigate. At 3:44 PM, MI began operating normally. The expert arrived at 3:49 PM and investigated but never saw a problem.

At 4:19 PM, Operations rested LRF5. This halted MiniBooNe beam.

A Pbar expert began reverse proton studies at 4:41 PM. Operator woke LRF5 for this injection of protons. MI experts also began a study period.

Tuesday May 27
The midnight shift began with Operations monitoring store 2614, with Pbar experts conducting studies, and with MiniBooNe taking beam, but at a reduced intensity,

At 12:28 AM, the Booster RF cavities started sparking. Operators investigated and turned the system off. Operations called in an expert who arrived at 2 AM.

The Booster expert discovered a RF phase problem at 4:11 AM. He turned this particular program off and enabled the injection capture program. He said that Operations could operate like this for studies. Repairs will begin after the planned power outage later this morning.

At 5 AM, Operations terminated store 2614. Safety techs began safety system tests.

Operations turned off the Linac at 5:30 AM and the TeV at 5:40 AM; all other areas followed.

The planned power outage occurred at 7 AM. The power came back eighteen minutes later. This power outage was the first step in the ComEd replacement of a bad high voltage power pole. This work will take three days. The Fermilab accelerator complex will conduct three days of maintenance.

By 8:30 AM, many areas were already in access.

Indications of a TeV LCW leak, at 9:31 AM, brought operators and water group techs together to search the TeV tunnels and buildings. The LCW leak was found near the D-zero (D0) collision hall. It was a ruptured LCW hose.

Wednesday May 28
The midnight shift began with the accelerator complex in shutdown, and with Operations monitoring systems.

At 1:10 AM, operators noticed a rise in cryo system temperature at TeV sector F4. After talking with cryo experts, operators successfully restarted the cooldown.

Today's Plan
Shutdown continues. The ComEd pole installation and removal is a little behind schedule, but the shutdown is still expected to end on Friday.

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/28/2003   email Fermilab